### v4

# Conformalized Strategy-Proof Auctions

Abstract
Auctions are key for maximizing sellers’ revenue and ensuring truthful bidding among buyers. Recently, an approach known as differentiable economics based on machine learning (ML) has shown promise in learning powerful auction mechanisms for multiple items and participants. However, this approach has no guarantee of strategy-proofness at test time. Strategy-proofness is crucial as it ensures that buyers are incentivized to bid their true valuations, leading to optimal and fair auction outcomes without the risk of manipulation. In this work, we propose a formulation of statistical strategy-proofness auction mechanism, ensuring that the probability of regret exceeding a predefined threshold is strictly controlled. Building upon conformal prediction techniques, we develop an auction acceptance rule that leverages regret predictions to guarantee that the data-driven auction mechanism meets the statistical strategy-proofness requirement with high probability. Our approach represents a practical middle-ground between two extremes: forcing zero-regret at the cost of significant revenue loss, and naively using ML to construct auctions with the hope of attaining low regret at test time. Numerical experiments demonstrate the necessity of the proposed method, the validity of our theoretical result, and its applicability.
1 Introduction
Auction design represents a fundamental component of economic theory, holding significant practical relevance in various industries and the public sector for organizing the sale of products and services [Milgrom, 2004, Krishna, 2009]. Auctions serve as a crucial mechanism for allocating scarce goods and generate billions in revenue across domains such as online advertising, sourcing, and more [Cramton, 1997, Sandholm, 2007, Leyton-Brown et al., 2017, Milgrom, 2017, Roth, 2018]. Notable examples include sponsored search auctions by search engines such as Google, and auctions on platforms such as eBay [Edelman et al., 2007]. Within the standard independent private valuations model, each bidder possesses a valuation function—the true values they are willing to pay—over subsets of items, independently drawn from possibly distinct distributions. The auctioneer is presumed to have knowledge of these value distributions, a crucial factor in auction design. However, a primary challenge arises from the private nature of valuations, as bidders might not disclose their true valuations, complicating the auction process.
1.1 Data-Driven Auctions
The pursuit of optimal auction mechanisms has been a subject of great interest to economics for many years. Myerson auction’s [Myerson, 1981] stands out as the optimal mechanism for auctioning a single item to several bidders. In contexts where multiple items are auctioned to a single bidder, the concept of optimality has been largely delineated [Manelli and Vincent, 2006, Pavlov, 2011, Daskalakis et al., 2015, Kash and Frongillo, 2016]. However, when it comes to auctioning multiple items to various bidders, optimal mechanisms have been identified only in simplified scenarios, such as those discussed by Yao [2017]. Indeed, the existence of an optimal mechanism in the multi-bidder multi-items setting remains an open question even in simple scenario such as two or more bidders and several items.
The predictive power of deep neural network (DNN) models, coupled with the quest for analytical outcomes, drive the research by Dütting et al. [2019] to a data-driven solution; see also [Dütting et al., 2024]. While this data-driven approach is not guaranteed to maximize revenue or be entirely strategy-proof, empirical evidence shows that the resulting deep-learning solutions are close to recapturing the outcome of well-known optimal auction designs, when they exist.
1.2 The Lack of Strategy-proofness in Data-Driven Auctions
Strategy-proofness relates closely to the notion of regret in auctions, a crucial metric that indicates whether the outcomes of a given auction mechanism are strategy-proof. Specifically, a mechanism with low regret ensures that bidders have little incentive to deviate from their true valuations. In other words, a low-regret mechanism encourages truthful bidding, which is a cornerstone of strategy-proofness [Dütting et al., 2015, Milgrom, 2004, Myerson, 1981]. As such, the importance of low regret in auctions cannot be overstated. In a practical sense, low regret translates to higher revenue for the auctioneer and more efficient allocation of resources among bidders. Moreover, it fosters trust in the auction mechanism, as participants can be confident that truthful bidding is in their best interest [Krishna, 2009].
A critical issue in deploying data-driven mechanisms for auctions, particularly those based on DNNs, is the lack of statistical guarantees during test time for new auctions. While these models are trained to minimize regret on average, leading to low-regret outcomes in many cases, they occasionally produce predictions for allocations and payments that grossly violate the low-regret requirement at test time. This limitation is visualized in Figure 1, illustrating that a well-known data-driven auction mechanism—RegretNet by Dütting et al. [2019]—fails to attain low-regret levels. While the average regret across many scenarios tends to be low, the presence of high regret values in the tail of the distribution is evident in the red zone in Figure 1. We stress this point in Figure 2, which shows an increase in the maximum regret obtained at test time as the number of auctions evaluated grows, suggesting that the model’s reliability diminishes as it encounters a wider variety of auction contexts.
Figure 1: Empirical distribution of actual regret values of RegretNet model, evaluated on 10,000 test auctions; the auction setup is $2\times 3$. For additional insights into the experimental configuration of this model, please refer to Section 4.
Figure 2: Maximum regret value obtained by RegretNet model as a function of the number of test auctions. The auction setup is $2\times 3$.
The discussion above reveals a significant limitation of data-driven auction design: there is no assurance that the regret will remain low with any level of confidence once the model is deployed in real-world scenarios. This undesired behaviour could result in substantial revenue losses for the auctioneer and unfair outcomes for the bidders, undermining the trust and efficiency of the auction system. The primary goal of this paper is to tackle this challenge by selectively accepting auctions for which the unknown test-time regret is unlikely to exceed a desired maximum regret level. However, before presenting our proposal, we first pause to review related work.
1.3 Related Work
Given the theoretical and practical importance of strategy-proofness, there has been growing activity in enhancing the reliability of data-driven auctions. Curry et al. [2020] modify the RegretNet model [Dütting et al., 2019] to enable certifiable strategy-proofness using mixed integer programming. However, the certification procedure imposes strict constraints both on the architecture design and the number of trainable parameters—as it involves solving a large mixed integer programming problem—limiting the model’s performance. By contrast, as we will see, the framework we propose in this paper can be applied to virtually any auction model, regardless of its complexity and architectural design.
A different line of work by Peri et al. [2021], Rahme et al. [2021b, a], Ivanov et al. [2022] offers novel learning approaches and architectures to formulate actions with low average regret. This stands in striking contrast with our method, which achieves a high probability bound on the worst-case, maximal regret obtained at test time.
Another line of work for forming reliable data-driven auctions builds on the parameterization of allocation menus that are guaranteed to be strategy-proof. This area was recently advanced by Curry et al. [2023] and Duan et al. [2023], who optimize the parameters of affine maximizer auctions (AMA) [Roberts, 1979]. While AMA-based approaches are strategy-proof, they often yield lower revenue than standard DNN-based methods. Indeed, in Section 4, we show that the approach by [Duan et al., 2023] results in a revenue loss of about 45% compared to our method. This emphasizes the hardness of attaining exact strategy-proof auctions while constructing data-driven auctions with high revenue.
1.4 Overview and Contributions
The challenge of forming zero-regret, strategy-proof auctions with high revenue drives us to propose a new, statistical approach for constructing reliable data-driven auctions. Our key contribution is the introduction of an auction acceptance rule that selectively rejects predicted auctions whose regret exceeds a pre-defined threshold. In turn, the advantage of our approach is that it guarantees, with high probability (e.g., 99%), that accepted auctions at test time do not exceed a pre-defined max regret value (e.g., of level 0.05).
Our approach consists of three key components. First, we present a novel, data-driven regret model to estimate the regret of an out-of-sample auction. We then build on conformal prediction to form a rigorous acceptance rule, ensuring that only auctions meeting the desired strategy-proofness level are accepted. Under the i.i.d. assumption, we prove that our approach controls the maximum regret level at test time, for any complex auction model, any unknown bidding distribution profiles, and regardless of the accuracy of the regret estimation model used to form the acceptance rule.
As such, our work introduces a new practical paradigm for balancing the trade-off between strategy-proofness and revenue discussed above. Our approach can be seen as a synthesis of two extremes: (i) enforcing zero regret at the expense of significant revenue loss, and (ii) naively using machine learning to design auctions with the hope of achieving low regret at test time. Indeed, our experiments show that the accepted auctions obtained by our approach are of low regret while having revenue nearly the same as the baseline auction model by Dütting et al. [2019].
2 Background
2.1 Problem Setup
We consider an auction environment, comprising of a set of $n$ bidders, denoted by $[n]=\{1,\ldots,n\}$, and a set of $m$ items, denoted by $[m]=\{1,\ldots,m\}$. Each bidder $i\in[n]$ possesses a private valuation $v^{i}\in\mathbb{R}^{m}$ which maps item sets $S\subseteq[m]$ to their value $v^{i}(S)\in\mathbb{R}_{\geq 0}$ for the bidder. The valuation matrix $v=[v^{1},\ldots,v^{n}]$ is assumed to be drawn from a distribution $P$.
In an auction, participants submit bids that may not necessarily represent their true valuations. The auction mechanism, denoted as $f(b)=[g(b),p(b)],$ comprises an allocation rule $g(b)$ and a payment rule $p(b)$. Specifically, the allocation rule is defined as $g:\mathbb{R}_{\geq 0}^{n\times m}\to[0,1]^{n\times m},$ determining how the $m$ items are assigned to the $n$ bidders. The payment rule is given by $p:\mathbb{R}_{\geq 0}^{n\times m}\to\mathbb{R}_{\geq 0}^{n\times m},$ specifying the amount each bidder must pay. Here, the bid profile of all bidders is represented as $b\in\mathbb{R}_{\geq 0}^{n\times m}$. The primary goal of the auction mechanism is to allocate the items efficiently while setting payments, often with the objective of maximizing revenue.
To effectively frame strategy-proof auctions as a learning task, we shall design a DNN trained to minimize the regret while maximizing the bidders’ utility. To this end, we denote the trained auction mechanism by ${f}_{\hat{\theta}}(b)=[{g}_{\hat{\theta}}(b),{p}_{\hat{\theta}}(b)]$, where the vector $\theta$ represents trainable parameters. Here, the ‘hat’ symbol over the parameters $\hat{\theta}$ indicates that these are fixed, i.e., the model is pre-trained. Using the notations from [Dütting et al., 2019], the utility of bidder $i$, denoted by $u^{i}_{\hat{\theta}}(v^{i};b)$, is defined as the difference between their valuation of the allocated items and the payment made, i.e.,
$u_{\hat{\theta}}^{i}(v^{i};b)=v^{i}({g}^{i}_{\hat{\theta}}(b))-{p}^{i}_{\hat{\theta}}(b).$ (1)
Above, $g^{i}_{\hat{\theta}}(b)\in\mathbb{R}$ and $p^{i}_{\hat{\theta}}(b)\in\mathbb{R}$ represent the allocation and payment for bidder $i$ given the bid profile $b$.
Using this utility function, the regret of bidder $i$ is defined as the difference in utility between bidding truthfully and bidding to maximize utility. Let $v^{-i}$ denote the valuation profile excluding bidder $i$. Then, the regret function of the $i$-th bidder is given by
$\operatorname{\textbf{rgt}}_{\hat{\theta}}^{i}(v;b)=\max_{b^{i}\in b}u_{\hat{\theta}}^{i}(v^{i};(b^{i},v^{-i}))-u_{\hat{\theta}}^{i}(v^{i};(v^{i},v^{-i})).$ (2)
Finally, the formulation of the learning task of an auction setup with valuations drawn from $P$ is as follows:
$\min_{\theta\in\mathbb{R}^{d}}$ $\quad\underset{v\sim P}{\mathbb{E}}\left[-\sum_{i\in N}p_{\theta}^{i}(v)\right]$ [REVENUE] (3)
s.t. $\operatorname{\textbf{rgt}}_{\theta}^{i}(v;b)=0,\hskip 7.39755pt\ \forall i\in[n]$  [REGRET]
$u_{\theta}^{i}(v^{i};b)\geq 0,\hskip 15.6491pt\forall i\in[n]$  [IR]
[…]
In plain words, the goal of this optimization problem is to maximize revenue while minimizing regret, promoting incentive rationality (IR) and a feasible mechanism.
2.2 Conformal Prediction
We now present the conformal prediction method, which we employ for controlling max regret levels within data-driven auction mechanisms.
Conformal prediction [Papadopoulos et al., 2002, Vovk et al., 2005, Lei and Wasserman, 2014] offers a generic approach, applicable to any predictive model, for quantifying prediction uncertainty. In recent years, conformal prediction has increasingly gained traction across various domains, demonstrating its broad utility and effectiveness in practical settings [Angelopoulos and Bates, 2023]. This method utilizes a subset of holdout calibration data to quantify the prediction error of the model. These errors are then used to construct a range of plausible predictions—a prediction interval—that covers the unknown test outcome with high probability.
In what follows, we describe in more detail how to construct prediction intervals for regression problems, as this task is tightly connected to our goal of reliably predicting unknown regret values at test time. Let $x\in\mathcal{X}$ represent the input features to the model, $y\in\mathcal{Y}$ denote the true output, and ${f}_{\hat{\theta}}(x)$ be the predicted output generated by the model ${f}_{\hat{\theta}}$ parametrized by ${\hat{\theta}}$. The process of conformal prediction for general inputs and outputs can be outlined as follows. First, we apply ${f}_{\hat{\theta}}$ on the labeled calibration data $\mathcal{D}_{\textup{cal}}=\{(x_{k},y_{k})\}_{k=1}^{K}$ to construct a heuristic notion of prediction error for each calibration point $(x_{k},y_{k})$. We refer to this notion of error as a nonconformity score, denoted by $s(x,y)\in\mathbb{R}$. The next step involves calculating $\hat{q}_{\alpha}$, the $(1-\alpha)(1+1/K)$ empirical quantile of the calibration scores $s(x_{k},y_{k})$, $k=1,\dots,K$. This quantile is used to construct a prediction interval for a new test point $x_{\textup{test}}$, which includes all the output values $\tilde{y}\in\mathcal{Y}$ whose corresponding nonconformity scores fall below the quantile $\hat{q}_{\alpha}$, i.e.
$C(x_{\textup{test}})=\{\tilde{y}\in\mathcal{Y}:s(x_{\textup{test}},\tilde{y})\leq\hat{q}_{\alpha}\}.$ (4)
The key property of $C(x_{\textup{test}})$ is that it is guaranteed to cover the unknown test label $y_{\textup{test}}$ at the desired level of coverage $1-\alpha$, as formally stated in Theorem 1.
Theorem 1 (Conformal Coverage Guarantee; Vovk, Gammerman, and Saunders 1999).
Assume that the calibration set ${\mathcal{D}_{\textup{cal}}=\{x_{i},y_{i}\}}_{k=1}^{K}$ and the test point $(x_{\textup{test}},y_{\textup{test}})$ are i.i.d. samples from $P_{x,y}$. Then, the following coverage guarantee holds for $\mathcal{C}(x_{\textup{test}})$ defined in (4):
$\mathbb{P}\Big{[}y_{\textup{test}}\in\mathcal{C}(x_{\rm test})\Big{]}\geq 1-\alpha.$ (5)
Notably, the standard conformal prediction approach cannot be applied “as is” in auction settings. This is because the regret value, which we aim to control, depends on the user’s bids $b$, valuations $v$, and the allocation and payment determined by the auction model ${f}_{\hat{\theta}}$. In other words, the “label” $y$ is now a complex function of both the model and the data, requiring a customized adaptation of conformal prediction to fit this auction context—this is the focus of our work.
3 Proposed Method
In this section, we present our approach to forming data-driven auctions with controlled maximal regret at test time. Given a test bidding matrix $b_{\textup{test}}$ and an unknown valuation matrix $v_{\textup{test}}$, we define the maximal regret as follows:
$\operatorname{\textbf{rgt}}_{\hat{\theta}}^{\textup{max}}(v_{\textup{test}};b_{\textup{test}})=\max_{i\in[n]}\ \operatorname{\textbf{rgt}}^{i}_{\hat{\theta}}(v_{\textup{test}};b_{\textup{test}}),$ (6)
where $\operatorname{\textbf{rgt}}^{i}_{\hat{\theta}}$ is defined in (2).
We also define the calibration data set $\mathcal{D}_{\textup{cal}}=\{(b_{k},v_{k})\}_{k=1}^{K}$, which contains holdout pairs of bidding $b_{k}\in\mathbb{R}^{n\times m}$ and valuation $v_{k}\in\mathbb{R}^{n\times m}$ matrices that correspond to the $k$-th instance. Throughout this paper, we assume that the calibration samples $(b_{k},v_{k}),k=1,\ldots,K$ and the test sample $(b_{\textup{test}},v_{\textup{test}})$ are drawn i.i.d. from some unknown distribution $P_{b,v}$.
Our approach consists of two key components. First, we develop a data-driven regret model, whose goal is to predict the unknown maximum regret of a new auction. Next, we show how to harness this regret estimator to construct an auction acceptance rule that accepts test actions that satisfy the following definition.
Definition 1 (Statistically Strategy-Proof Auctions).
An auction mechanism ${f}_{\hat{\theta}}(b_{\textup{test}})=[{g}_{\hat{\theta}}(b_{\textup{test}}),{p}_{\hat{\theta}}(b_{\textup{test}})]$ is said to be statistically strategy-proof at a user-specified confidence level $(1-\alpha)\in[1,0)$ and maximal regret level $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$ if
$\mathbb{P}\left[\textup{{rgt}}^{\textup{max}}_{\hat{\theta}}(v_{\textup{test}};b_{\textup{test}})>\operatorname{\ell_{\operatorname{\textbf{rgt}}}}\right]\leq\alpha,$ (7)
where the probability is taken over calibration and test points.
In words, an auction outcome that satisfies Definition 1 is guaranteed to have low maximum regret with high probability, where the maximal regret value $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$ and the confidence level $1-\alpha$ are pre-defined by the auctioneer. Since we can control the maximal regret value with high probability, we refer to it as a statistically strategy-proof auction mechanism. Notably, this statistical notion of strategy-proofness does not imply zero regret at test time.
3.1 Regret Estimation
Since the regret of a new auction is unknown, we should introduce a regret estimation model. In its most general form, we formulate it as a neural network ${r}_{{\hat{\theta}}_{r}}(b)$, parameterized by ${\hat{\theta}}_{r}$; this model predicts
$\operatorname{\textbf{rgt}}_{\hat{\theta}}(v;b)=[\operatorname{\textbf{rgt}}^{1}_{\hat{\theta}}(v;b),\operatorname{\textbf{rgt}}^{2}_{\hat{\theta}}(v;b),\ldots,\operatorname{\textbf{rgt}}^{n}_{\hat{\theta}}(v;b)],$
the outcome regret for a given auction. To fit the regret model on the training set $\mathcal{D}_{\textup{train}}$, we minimize the following objective function:
$\min_{\theta_{r}\in\mathbb{R}^{d}}\sum_{(b,v)\in\mathcal{D}_{\textup{train}}}\|r_{\theta_{r}}(b)-\operatorname{\textbf{rgt}}_{\hat{\theta}}(v;b)\|_{1}.$ (8)
We propose two different approaches to design the regret estimation model ${r}_{{\hat{\theta}}_{r}}(b)$. In the shared backbone setup illustrated in Figure 3, the regret model is integrated into the existing allocation and payment backbone, providing a unified structure. The advantage of this approach is that it leverages the backbone features of the auction model to better predict the regret. However, this setup requires access to the auction model parameters and modifies the training procedure. In the second black-box approach, the auction model is treated as a black-box, avoiding the need to access the auction model’s internal parameters. This design choice is visualized in Figure 4, where the regret estimation model is a standalone neural network.
Figure 3: Shared backbone regret estimation.
Figure 4: Black-box auction model regret estimation.
3.2 Model Calibration
In this section, we introduce a data-driven auction acceptance rule that accepts auctions with regret outcomes below $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$—the maximum regret level specified by the auctioneer. A naive approach would be to accept a new auction ${f}_{\hat{\theta}}(b_{\text{test}})$ if it satisfies ${r}_{\hat{\theta}_{r}}^{\textup{max}}(b_{\text{test}})\leq\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$, where ${r}_{\hat{\theta}_{r}}^{\textup{max}}(b_{\textup{test}})=\max_{i\in[n]}{r}_{\hat{\theta}_{r}}^{\textup{i}}(b_{\textup{test}})$. In plain words, the auction is accepted if its estimated maximal predicted regret falls below the threshold $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$. However, the key issue with this naive approach is that the regret estimator may be inaccurate, potentially leading to the acceptance of auctions whose actual regret $\operatorname{\textbf{rgt}}^{\textup{max}}_{\hat{\theta}}(v_{\text{test}};b_{\text{test}})$ exceeds $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$. This emphasizes the importance of the calibration procedure presented in this section, which rigorously bridges the gap between the estimated maximal regret ${r}^{\textup{max}}_{\hat{\theta}_{r}}(b_{\text{test}})$ and the actual, unknown maximal regret $\operatorname{\textbf{rgt}}^{\textup{max}}_{\hat{\theta}}(v_{\text{test}};b_{\text{test}})$.
In essence, we show how to form a corrected threshold on the output ${r}^{\textup{max}}_{\hat{\theta}_{r}}(b_{\text{test}})$, ensuring the regret of the auction outcome obtained from the auction model does not exceed the requested maximal regret level $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$ with probability $1-\alpha$, resulting in a calibrated auction acceptance rule function. Given the auction model ${f}_{\hat{\theta}}$, a calibration set $\mathcal{D}_{\textup{cal}}$, a requested regret level $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$, and a control level $\alpha$, we define the auction acceptance rule function $\hat{\mathcal{T}}_{\alpha}(b_{\textup{test}},\operatorname{\ell_{\operatorname{\textbf{rgt}}}})$ for a test bid vector $b_{\text{test}}$ as follows:
$\hat{\mathcal{T}}_{\alpha}(b_{\text{test}},\operatorname{\ell_{\operatorname{\textbf{rgt}}}})=\begin{cases}{f}_{\hat{\theta}}(b_{\text{test}}),&{r}^{\textup{max}}_{\hat{\theta}_{r}}(b_{\text{test}})\leq\operatorname{\ell_{\operatorname{\textbf{rgt}}}}-\hat{q}_{\alpha},\\ \;\;\;\emptyset,&\text{otherwise}.\end{cases}$ (9)
The function $\hat{\mathcal{T}}_{\alpha}(b_{\text{test}},\operatorname{\ell_{\operatorname{\textbf{rgt}}}})$ from above is determined by the estimated maximum regret ${r}^{\textup{max}}_{\hat{\theta}_{r}}(b_{\text{test}})$ and the calibration threshold $\hat{q}_{\alpha}$, which is derived using the calibration set $\mathcal{D}_{\textup{cal}}$. The output of the function $\hat{\mathcal{T}}_{\alpha}(b_{\text{test}},\operatorname{\ell_{\operatorname{\textbf{rgt}}}})$ is the prediction of the auction model ${f}_{\hat{\theta}}({b_{\text{test}}})$ if ${r}^{\textup{max}}_{\hat{\theta}_{r}}(b_{\text{test}})$ is lower than the difference between the requested regret level $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$, and the correction factor $\hat{q}_{\alpha}$ derived from our calibration method. Otherwise, the function $\hat{\mathcal{T}}_{\alpha}$ yields an empty set $\emptyset$, indicating that the predicted auction does not meet the desired regret level. This acceptance rule ensures that the auction mechanism only operates within the predefined regret bounds.
Algorithm 1 details the calibration process. The procedure starts by computing the regret scores $e_{k}=\operatorname{\textbf{rgt}}^{\textup{max}}_{\hat{\theta}}(v_{k};b_{k})-{r}^{\textup{max}}_{\hat{\theta}_{r}}(b_{k})$, defined as the difference between the actual and the estimated regrets for each bidding $b_{k}$ and valuation $v_{k}$ pair of the calibration dataset. Then, we compute the empirical quantile $\hat{q}_{\alpha}$ of the regret scores $e_{k}$, $k=1,\ldots,K$. This quantile forms our threshold, which is used to set the corrected cutoff in the acceptance rule $\hat{\mathcal{T}}_{\alpha}$ from (9). That is, $\hat{q}_{\alpha}$ rigorously accounts for the prediction error of the regret estimation model ${r}_{\hat{\theta}_{r}}$.
Algorithm 1 Regret Calibration
1:Input: ${f}_{\hat{\theta}}$ – pre-trained auction model; $r_{\hat{\theta}_{r}}$ – pre-trained regret model; $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$ – max regret level; $(1-\alpha)$ – requested confidence level; $\mathcal{D}_{\textup{cal}}$ – holdout calibration set.
2:Initialize an empty list $R$ to store regret scores.
3:for  $(b_{k},v_{k})$ in $\mathcal{D}_{\textup{cal}}$ do
4:     Compute the estimated maximum regret ${r}_{\hat{\theta}_{r}}^{\textup{max}}(b_{k})$.
5:     Compute the actual maximum regret $\operatorname{\textbf{rgt}}_{\hat{\theta}}^{\textup{max}}(v_{k};b_{k})$.
6:     Calculate the score $e_{k}=\operatorname{\textbf{rgt}}_{\hat{\theta}}^{\textup{max}}(v_{k};b_{k})-{r}_{\hat{\theta}_{r}}^{\textup{max}}(b_{k})$.
7:     Append $e_{k}$ to $R$.
8:end for
9:Compute $\hat{q}_{\alpha}:=$ the $\lceil(1-\alpha)(|\mathcal{D}_{\textup{cal}}|+1)\rceil$-th smallest element in $\{e_{i}\in R\}\cup\{\infty\}$.
10:return Calibrated acceptance rule function $\hat{\mathcal{T}}_{\alpha}(b,\operatorname{\ell_{\operatorname{\textbf{rgt}}}})$, defined in (9) and implemented with $\hat{q}_{\alpha}$.
Table 1: Summary of the average revenue and empirical regret, with their standard deviations in parentheses, for both RegretNet and ConfRegretNet across $2\times 2$, $2\times 3$, and $3\times 5$ auction settings. The test and the calibration sets are of size 10,000 examples. The ‘Requested Max Regret Level‘ column indicates the requested regret used in ConfRegretNet to control the regret at $99\%$ confidence ($\alpha=0.01$). The ‘Accepted Auctions‘ column shows the percentage of accepted auctions by ConfRegretNet based on the regret threshold. the ‘Max Regret‘ column refers the max regret obtained at test time.
Auction Model Auctio Setup Bidders × Items Revenue Empirical Regret Requested Max Regret Max Regret Acceptance Auctions
RegretNet $2\times 2$ 0.93 (0.33) 0.007 (0.006) - 0.039 -
ConfRegretNet [ours] $2\times 2$ 0.92 (0.32) 0.004 (0.003) 0.025 0.020 89.23%
RegretNet $2\times 3$ 1.41 (0.38) 0.011 (0.014) - 0.164 -
ConfRegretNet [ours] $2\times 3$ 1.38 (0.39) 0.007 (0.005) 0.035 0.032 94.4%
RegretNet $3\times 5$ 2.93 (0.40) 0.023 (0.012) - 0.121 -
ConfRegretNet [ours] $3\times 5$ 2.67 (0.38) 0.009 (0.007) 0.055 0.055 93.4%
We now prove that any test auction accepted by the rule $\hat{\mathcal{T}}_{\alpha}$ satisfies Definition 1. Simply put, this result guarentees that by setting $\alpha=0.01$ and specifying a maximal regret level of $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}=0.025$, the actual regret of new auctions accepted by $\hat{\mathcal{T}}_{\alpha}$ will not exceed 0.025 with probability 99%.
Proposition 1.
Suppose the calibration set $\mathcal{D}_{\textup{cal}}=\{(b_{k},v_{k})\}_{k=1}^{K}$ and a new auction $(b_{\textup{test}},v_{\textup{test}})$ are i.i.d samples from $P_{b,v}$. Let $\hat{\mathcal{T}}_{\alpha}$ be the acceptance rule of a given auction model ${f}_{\hat{\theta}}$ and regret estimator $r_{\hat{\theta}_{r}}$, obtained from Algorithm 1. Then, the auction ${f}_{{\hat{\theta}}}(b_{\textup{test}})$ accepted by the rule $\hat{\mathcal{T}}_{\alpha}$ in (9) satisfies Definition 1 with regret level $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$ and confidence level $1-\alpha$.
Proof.
To lighten notations, denote by $r$ the actual maximum regret $\operatorname{\textbf{rgt}}^{\textup{max}}_{\hat{\theta}}(v_{\textup{test}};b_{\textup{test}})$ and by $\hat{r}$ the predicted maximum regret ${r}_{\hat{\theta}_{r}}^{\textup{max}}(b_{\textup{test}})$. With these notations in place, the probability of interest is as follows:
$\mathbb{P}\left[(r>\operatorname{\ell_{\operatorname{\textbf{rgt}}}})\>\cap\>(\hat{\mathcal{T}}_{\alpha}(b_{\textup{test}},\operatorname{\ell_{\operatorname{\textbf{rgt}}}})\neq\emptyset)\right]$
$=\>\mathbb{P}\left[(r>\operatorname{\ell_{\operatorname{\textbf{rgt}}}})\>\cap\>(\hat{r}<\operatorname{\ell_{\operatorname{\textbf{rgt}}}}-\hat{q}_{\alpha})\right]$ by definition of  $\hat{\mathcal{T}}_{\alpha}$
$=\>\mathbb{P}\left[(r>\operatorname{\ell_{\operatorname{\textbf{rgt}}}})\>\cap\>(\operatorname{\ell_{\operatorname{\textbf{rgt}}}}>\hat{r}+\hat{q}_{\alpha})\right]$
$=\>\mathbb{P}\left[r>\hat{r}+\hat{q}_{\alpha}\right]$ intersecting events
$=\>\mathbb{P}\left[r-\hat{r}>\hat{q}_{\alpha}\right]$
= $=\>1-\mathbb{P}\left[r-\hat{r}\leq\hat{q}_{\alpha}\right]$ complement event
 $\leq\>1-(1-\alpha)$
$=\>\alpha,$
where the last inequality holds due [Romano et al., 2019, Lemma 2]. ∎
4 Experiments
In this section, we empirically evaluate our proposal on auctions with configurations of (bidders $\times$ items) sets of size $2\times 2$, $2\times 3$, and $3\times 5$. The training process is based on the RegretNet implementation by Dütting et al. [2019] (PyTorch version by Stein et al. [2023]), which is used as the data-driven auction model $f_{\hat{\theta}}$. Throughout the experiments, we refer to our conformal method as ConfRegretNet. That is, we apply the acceptance rule $\hat{\mathcal{T}}_{\alpha}$ to the output of the auction model ${f}_{\hat{\theta}}(b)=[{g}_{\hat{\theta}}(b),{p}_{\hat{\theta}}(b)]$. We fit our regret estimation model ${r}_{\hat{\theta}_{r}}$ within the shared backbone architecture setup, illustrated in 3. For further implementation details, please refer to Section A in the Appendix.
The performance of our ConfRegretNet method is evaluated in terms of revenue, empirical regret, and maximum regret. Table 1 summarizes the results for different auction configurations. Following that table, we can see our method achieves competitive revenue compared to RegretNet across different auction settings while controlling the requested levels of maximum regret. Observe how the maximum regret values obtained by RegretNet are uncontrolled.
(a) Auction setup 2x2
(b) Auction setup 2x3
(c) Auction setup 3x5
Figure 5: The distribution of regret values corresponding to RegretNet and ConfRegretNet across different auction settings: $2\times 2$, $2\times 3$, and $3\times 5$. The test and the calibration sets are of size 1,000 examples each. The red dashed line indicates the requested regret level $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$ of each setup.
Figure 6: Scatter plot of the actual regret $\operatorname{\textbf{rgt}}_{\hat{\theta}}(v;b)$ and the predicted regret ${r}_{\hat{\theta}_{r}}(b)$ of ConfRegretNet, evaluated on test data of size $1,000$; the auction setup is $2\times 2$.
Table 1 also shows that our acceptance rule tends to accept most of the auctions at test time. Further, Figure 5 illustrates that (i) the auctions that are rejected by our method are indeed those that exceed the specified regret level; and (ii) the regret values of the accepted auctions fall below the maximal regret level $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$. This implies that our approach not only controls the regret level but also tends to accept auctions that should be accepted.
Figure 6 presents the actual regret values versus the predicted regret values from our regret estimation model ${r}_{\hat{\theta}_{r}}$. The strong correlation observed between them suggests that the model predicts regret with reasonable accuracy. This aligns with the behavior seen in Figure 5, where our method consistently rejects auctions with high regret.
Table 2 summarizes the performance of various classical, transformer-based, and deep learning-based auction mechanisms in the $2\times 2$ setting. Among the strictly strategy-proof methods, AMenuNet achieves a higher revenue (0.43) than the classical VCG (0.28) and Item-Myerson (0.42). However, better revenues are achieved by the neural network models—RegretFormer (0.94), RegretNet (0.93), ALGNet (0.88), and EquivariantNet (0.87)—all of which surpass the classical approaches. Notably, ConfRegretNet attains revenue (0.92) comparable to these high-performing models while offering a statistical strategy-proofness guarantee. This demonstrates the practical trade-off between guaranteeing zero-regret and maximizing revenue, highlighting ConfRegretNet as a compelling option for real-world auction applications where both factors matter.
Table 2: Comparison of average revenue and strategy-proofness of various auction models in the $2\times 2$ auction setting. The table includes our ConfRegretNet compared to classical mechanisms (VCG, Item-Myerson), transformer-based models (AMenuNet, RegretFormer), and other deep learning-based approaches (RegretNet, EquivariantNet, ALGNet). Strategy-proofness is indicated as either Yes, No, or Statistically. The results presented in this table are as reported by the authors for the same setup.
Auction Model Revenue Strategy-Proof
VCG [Vickrey, 1961] 0.28 Yes
Item-Myerson [Myerson, 1981] 0.42 Yes
AMenuNet [Duan et al., 2023] 0.43 Yes
RegretNet [Dütting et al., 2019] 0.93 No
EquivariantNet [Rahme et al., 2021a] 0.87 No
ALGNet [Rahme et al., 2021b] 0.88 No
RegretFormer [Ivanov et al., 2022] 0.94 No
ConfRegretNet [ours] 0.92 Statistically
To demonstrate the flexibility of our approach, we now conduct experiments in which we treat the original RegretNet model as a black-box and form an acceptance rule by training a separate regret model. The results, summarized in Table 3, show the validity of our approach and our ability to perform well even without having access to the auction model’s parameters.
Table 3: Summary of the average revenue and empirical regret (with standard deviations in parentheses) for RegretNet, ConfRegretNet-shared (3), and ConfRegretNet-separate (4) in the black-box settings described in Section 3.1. Each model’s requested and actual maximum regret is reported, along with the percentage of auctions accepted under the requested regret threshold. Test and calibration sets each contain 10,000 examples.
Auction Model Auction Setup Bidders × Items Revenue Empirical Regret Requested Max Regret Max Regret Acceptance Auctions
RegretNet $2\times 2$ 0.93 (0.33) 0.007 (0.006) - 0.039 -
ConfRegretNet-shared (3) $2\times 2$ 0.92 (0.32) 0.004 (0.003) 0.025 0.020 89.23%
ConfRegretNet-separate (4) $2\times 2$ 0.92 (0.31) 0.004 (0.005) 0.025 0.023 93.7%
Lastly, Figure 7 presents the effect of the maximal regret level $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$ on the number of auctions rejected by $\hat{\mathcal{T}}_{\alpha}$. Following that figure, we can see that as the auctioneer enforces a lower max regret level $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$, the number of rejected auctions increases as they fail to meet this stricter criterion.
Figure 7: The effect of the maximal allowed regret level $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$ on the number of auctions rejected by the calibrated acceptance rule $\hat{\mathcal{T}}_{\alpha}(b,\operatorname{\ell_{\operatorname{\textbf{rgt}}}})$. Results are evaluated on a test set of size 400; the auction setup is $2\times 2$.
Overall, the results validate the effectiveness of our approach in designing auctions that balance revenue maximization with approximate strategy-proofness.
5 Discussion
In this paper, we introduced the notion of statistical strategy-proof auctions, and formulated a data-driven acceptance rule that selectively accepts auctions, at test time, whose regret is guaranteed to fall below a user-specified maximum regret value with high probability. The formulation of our data-driven acceptance rule is achieved based on predictions obtained by our novel regret estimation model, which is then calibrated by building on ideas from the conformal prediction literature.
Numerical experiments showed that while the average regret of the vanilla RegretNet model is low, its maximum regret can be relatively high. By applying our proposed acceptance rule, the regret of accepted auctions consistently remained within the prescribed maximum regret level. Notably, the results demonstrated the effectiveness of our approach in filtering out high-regret auctions while retaining those with low regret. The experiments also demonstrated that our approach achieved notably higher revenue compared to baseline methods that enforce strictly strategy-proof auctions with zero regret, highlighting the practical relevance of our statistical approach.
Limitations and future directions
The usefulness of our acceptance rule is affected by the predictive power of the underlying regret estimation model. If this model has poor predictive capabilities, our algorithm may make more rejections to ensure that the maximal regret will be controlled. Future research could benefit from exploring various learning schemes to predict the unknown regret more effectively.
A natural limitation of our method is the i.i.d. assumption we make, which might be violated in practice. This may happen due to changes in the bidding and valuation distribution profiles, or due to imperfect calibration data that include noisy or inaccurate valuations, for instance. As a future research direction, it would be important to quantify the effect of violation of this assumption on the resulting maximal regret level, possibly by taking inspiration from Barber et al. [2023]. It would be also intriguing to design calibration frameworks that are robust to such distribution shifts in the data.
Another worthwhile direction for future work is to extend our framework to simultaneously control the test-time revenue and regret. We believe this could be done by borrowing ideas from [Angelopoulos et al., 2021], which presents a general calibration framework for controlling multiple risks. Such an approach would enable the auctioneer to ensure, with high probability, that the calibrated auction mechanism maintains low regret while also keeping the revenue of the acceptance auctions above a specified threshold. This multi-objective control is particularly relevant for large-scale, multi-item auctions, where revenue could potentially fall dramatically due to a large number of items.
Acknowledgements.
Y.R. and R.M.L. were supported by the ISRAEL SCIENCE FOUNDATION (grant No. 729/21). Y.R. thanks the Career Advancement Fellowship, Technion. I.T.C received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement: 101077862, project name ALGOCONTRACT, PI: Inbal Talgam-Cohen).
Supplementary Material
Appendix A Architectural and Training Details
The architecture of our models includes ReLU activations for both the allocation and payment networks, as well as the regret estimation model if exists. To ensure that each item’s allocations sum to one, we apply a softmax activation to the allocation network. For the payment network, a sigmoid activation function is used to constrain the payments to a fraction between zero and one, enforcing individual rationality (IR).
Models are trained on 700,000 sample valuation profiles $b_{i}\sim U[0,1]$, divided into batches, using the Adam optimizer with an initial learning rate of $1e^{-3}$. The training process spans several epochs, with the exact number described in Table 4.
An additional term for regret minimization is incorporated into the training process, similar to the augmented Lagrangian method used in RegretNet. The weight for this regret term is periodically updated to adjust the emphasis on regret minimization in the loss function. Furthermore, the value of the parameter $\rho$, which influences the regret term, is also periodically updated during training.
The agents’ regret is optimized in an inner loop, with specific parameters detailed in the last three rows of Table 4. For a comprehensive understanding of the loss function derivation and further details on the training procedure, we refer the reader to the original RegretNet paper by Dütting et al. [2019].
Table 4: Hyperparameters. We provide training hyperparameters for all auction sizes we study.
Auction setup (Bidders x Items) 2$\times$2 2$\times$3 3$\times$5
Epochs 50 50 50
Train Batch Size 512 1024 2048
Initial Learning Rate 0.001 0.005 0.01
Number of Hidden Layers 5 5 5
Hidden Layer size 100 100 100
$\rho$ Update Period (Epochs) 2 2 2
Lagrange Weight Update Period (Iters) 100 100 100
Initial $\rho$ 1 1 1
$\rho$ Increment 1 5 8
Initial Lagrange Weight 5 5 5
Misreport Learning Rate (Training) 0.1 0.1 0.1
Misreport Iterations (Training) 25 25 25
Misreport Initializations (Training) 10 10 10
Appendix B Experiment setup and environment
The experiments were conducted on a high-performance computing setup consisting of an Ubuntu 20.04.6 LTS operating system, powered by 96 Intel(R) Xeon(R) Gold CPUs running at 2.40 GHz. The machine was equipped with 16 Nvidia A40 GPUs and 512 GB of RAM. The software environment was based on PyTorch 2.1, running under Python 3.11.5, which ensured compatibility with the (current) latest available libraries and frameworks.